# SyncTwin: Fast Digital Twin Construction and Synchronization for Safe Robotic Grasping

### v1

@@ 4.5 Ablation Studies
Table 2: Comparison of grasp success rates before and after geometry completion. Asset-based completion significantly improves performance across all objects by producing more accurate and safer grasp pose candidates, with the largest gains observed for partially occluded items such as the cup (with handle).
@@
Object Completion. We evaluate the effect of object completeness on grasp generation. When grasp poses are generated from single-view partial point clouds, the limited geometry often leads to incomplete or incorrect grasp configurations, for example. As shown in Figure 9, the cup handle is misinterpreted, resulting in unsafe or colliding grasps. In contrast, after SyncTwin aligns the observed object with the corresponding complete mesh from the asset library, the grasp generator produces denser, more accurate, and physically feasible grasp candidates. This demonstrates that asset-based completion is essential for reliable and collision-free grasp execution in the real world. As shown in Table 2, across 60 evaluation trials, grasp success rates increase substantially after geometry completion. The improvement is most pronounced for partially occluded or asymmetric objects, such as the cup and chips can, where asset-based completion provides the missing structure needed for reliable planning.
@@
4.6 Discussions
In our experiments, Stage I reconstruction occasionally fails when the object lacks a clear supporting surface or when the support plane is weakly connected to the object (e.g., a tall cup). In these cases, inaccurate camera extrinsics cause the 2D segmentation masks to be incorrectly projected into 3D, resulting in broken or incomplete object point clouds and missing geometry. Such extrinsic errors produced by learning-based reconstruction remain an open challenge.
5 Conclusion
We introduced SyncTwin, a digital-twin framework that unifies fast RGB-only 3D reconstruction with real-time scene synchronization for safe and robust grasping in dynamic, partially occluded environments. By leveraging VGGT-based reconstruction, segmentation-aware denoising, and memory-driven geometry completion, SyncTwin provides accurate object geometry and reliable grasp generation from limited visual input. By bridging the sim-to-real gap through consistent digital-twin updates, the system enables simulation-based planners to execute safe, collision-aware trajectories on real robots without retraining.
Our experiments demonstrate substantial improvements over existing baselines such as NVBlox, including higher obstacle-avoidance success rates, more stable behavior in a dynamic environment, and improved grasp performance under single-view occlusion. These results highlight the advantage of combining fast asset generation with persistent memory and real-time synchronization, enabling safe and accurate execution in real-world environments.
Looking forward, two extensions appear particularly promising. First, enabling online asset expansion by integrating Stage I construction into Stage II synchronization would allow the system to incrementally acquire new objects as they appear. Second, distributed or multi-GPU system designs that decouple perception, synchronization, and planning could further improve responsiveness under fast dynamic scenes. Consequently, these enable more adaptive, real-world–oriented digital-twin synchronization.
Supplementary Material
6 Denoising Algorithms Details
@@
$\theta_{i}=\arccos\!\left(1-\frac{2i}{K}\right),\qquad\phi_{i}=\pi(1+\sqrt{5})\,i,$ (4)
@@
$\mathbf{p}^{*}_{j}=\arg\max_{\mathbf{p}_{i}\in\mathcal{P}}\|\mathbf{p}_{i}-\mathbf{c}\|,\qquad\text{s.t. }\hat{\mathbf{v}}_{i}\cdot\mathbf{d}_{j}\geq\cos(\theta_{\text{tol}}).$ (12)
@@
7 Green Cube-Based Scale Estimation Algorithm
@@
Among several detection masks, the method yielding a reasonable number of points (100010005000050000) is selected to ensure tight coverage of the cube. The result is shown in Figure 11.
@@
The bounding box of detected green points is computed to estimate cube dimensions. A refined analysis using PCA is then applied to:
@@
Figure 11: Green cube–based scale estimation. The top image shows the whole reconstructed point cloud. The bottom illustrates the isolated cube points and the fitted red bounding box. By measuring the cube’s average edge length in VGGT-normalized coordinates and comparing it to the known real-world cube size (e.g., $0.1\,\mathrm{m}$), the scale factor between VGGT coordinates and real-world metric space is computed.
@@
8 3D Asset Visualization
@@
9 Tips for Practical Implementation
@@
1) VGGT Extrinsics Are Not Stored in the GLB File.
@@
3) Asset Centering for Consistent Pose Alignment.
@@
4) Distinguishing camera_optical from RealSense Physical Extrinsics.
@@
6) Using GraspGen with SyncTwin.
@@
7) NVBlox Filtering Range Adjustment.
@@
8) Synchronization Requirements for cuRobo.

### v2

@@ IV-F Ablation Studies
TABLE III: Comparison of grasp success rates before and after geometry completion using GraspGen. GraspGen is a state-of-the-art point-cloud-based grasp pose generator, and asset-based completion significantly improves its performance across all objects by providing more complete object geometry, resulting in more accurate and safer grasp pose candidates.
@@
Object Completion. We evaluate the effect of object geometric completeness on grasp generation. When grasp poses are generated from single-view partial point clouds, the limited geometry often leads to incomplete or incorrect grasp configurations. For example, as shown in Fig. 10, the handle of the cup is not perceived, resulting in unsafe or colliding grasps. In contrast, after SyncTwin aligns the observed object with the corresponding complete mesh from the asset library, the grasp generator produces denser, more accurate, and geometrically feasible grasp candidates. This demonstrates that asset-based completion is essential for reliable and collision-free grasp execution in the real world. As shown in Table III, across 60 evaluation trials, grasp success rates increase substantially after geometry completion. The improvement is most pronounced for partially occluded or asymmetric objects, such as the cup with a handle and the chips can, where asset-based completion provides the missing structure needed for reliable planning.
@@
IV-G Discussions
Our work provides a segmentation approach that covers most objects, while some challenging cases remain when applied to learning-based reconstructed point clouds. For instance, the supporting-plane denoising strategy may fail for tall or weakly supported objects such as wine glasses. To the best of our knowledge, this work represents the first attempt to address this problem, and many challenges still remain.
This work emphasizes a principled view of digital twins as real-time synchronization of environmental geometry, rather than merely visual or appearance-level simulation. Although several prior works are referred to as digital twins, they primarily focus on visual consistency and perform a single offline scene reconstruction, without explicitly addressing continuous state synchronization between the real world and simulation. As a result, such systems are fundamentally closer to visual simulation than to a digital twin with real-time synchronization capability. In contrast, SyncTwin is explicitly designed to continuously track and synchronize object geometry and pose, capturing dynamic changes in the environment and enabling faithful real-to-sim-to-real state consistency.
Moreover, the proposed digital twin synchronization also provides a foundation for safe real-robot reinforcement learning. By synchronizing real-world geometry into simulation, the system enables collision-aware training and execution, thereby reducing the risk associated with real-world exploration. This capability is particularly important for post-training and fine-tuning learning-based policies on physical robots.
V Conclusion
We introduced SyncTwin, a digital twin framework that unifies fast RGB-only 3D reconstruction with real-time scene synchronization for safe and robust execution in dynamic, partially occluded environments. By combining learning-based reconstruction, segmentation-aware denoising, and memory-driven geometry completion, SyncTwin provides accurate object geometry and reliable grasp generation from limited visual input, enabling simulation-based planners to execute safe, collision-aware trajectories on real robots without retraining.
Our experiments demonstrate substantial improvements over existing baselines such as NVBlox, including higher obstacle-avoidance success rates, more stable behavior in dynamic environments, and improved grasp performance under single-view occlusion. These results highlight the effectiveness of integrating fast asset generation with persistent memory and real-time synchronization for reliable real-world execution.
Looking forward, several extensions are promising. Integrating Stage I construction into Stage II synchronization would enable online asset expansion, allowing the system to incrementally acquire new objects. Extending SyncTwin beyond geometry-level synchronization to visual appearance consistency would support more realistic simulation and training of visuomotor policies. In addition, distributed or multi-GPU system designs that decouple perception, synchronization, and planning could improve responsiveness in highly dynamic scenes. Together, these directions enable adaptive and scalable digital twin synchronization for real-world robotic systems.
Appendix A Denoising Algorithms Details
@@
$\theta_{i}=\arccos\!\left(1-\frac{2i}{K}\right),\quad\phi_{i}=\pi(1+\sqrt{5})\,i,$ (4)
@@
Figure 11: Green cube-based scale estimation. The top image shows the whole reconstructed point cloud. The bottom illustrates the isolated cube points and the fitted red bounding box. By measuring the cube’s average edge length in VGGT-normalized coordinates and comparing it to the known real-world cube size (e.g., $0.1\,\mathrm{m}$), the scale factor between VGGT coordinates and real-world metric space is computed.
@@
$\mathbf{p}^{*}_{j}=\arg\max_{\mathbf{p}_{i}\in\mathcal{P}}\|\mathbf{p}_{i}-\mathbf{c}\|,\quad\text{s.t. }\hat{\mathbf{v}}_{i}\cdot\mathbf{d}_{j}\geq\cos(\theta_{\text{tol}}).$ (12)
@@
Appendix B Green Cube-Based Scale Estimation Algorithm
@@
Among several detection masks, the method yielding a reasonable number of points (100010005000050000) is selected to ensure tight coverage of the cube. The result is shown in Fig. 11.
@@
The bounding box of detected green points is computed to estimate the cube dimensions. A refined analysis using PCA is then applied to:
@@
Appendix C 3D Asset Visualization
We construct object assets using RGB images captured by a standard camera (e.g., a mobile phone), rather than relying on multiple depth sensors, for several practical reasons.
First, RGB cameras are low-cost and widely available, allowing asset acquisition to be performed easily using commodity devices such as smartphones. Second, our system already employs a single RealSense camera that is fixed in the environment for online perception and tracking.
Using this depth camera for asset construction would require moving it around the scene to capture multiple viewpoints, followed by repeated extrinsic calibration each time the camera is re-mounted. In practice, this calibration process is time-consuming and cumbersome, especially when assets need to be constructed for different environments. By contrast, using a handheld RGB camera avoids repeated extrinsic calibration entirely, enabling flexible multi-view data capture by simply walking around the scene. This design choice significantly reduces the overhead of asset acquisition and improves deployability in real-world settings.
Finally, while some prior works rely on fixed cameras and rotate the target object during reconstruction, this assumption does not hold in many realistic scenarios. For large objects or static environments such as tables, shelves, or cabinets, rotating the object is infeasible. Our RGB-based asset construction strategy naturally supports such settings by allowing viewpoint changes without requiring object manipulation.
@@
Appendix D Tips for Practical Implementation
@@
1) VGGT Extrinsics Are Not Stored in the GLB File
@@
3) Asset Centering for Consistent Pose Alignment
@@
4) Distinguishing camera_optical from RealSense Physical Extrinsics
@@
6) Using GraspGen with SyncTwin
@@
7) NVBlox Filtering Range Adjustment
@@
8) Synchronization Requirements for cuRobo